### v1

# A Survey on Class-Agnostic Counting: Advancements from Reference-Based to Open-World Text-Guided Approaches

Abstract
Object counting has recently shifted towards class-agnostic counting (CAC), which addresses the challenge of counting objects across arbitrary categories, tackling a critical need in versatile counting systems. While humans effortlessly identify and count objects from diverse categories without prior knowledge, most counting methods remain restricted to enumerating instances of known classes, requiring extensive labeled datasets for training, and struggling under open-vocabulary settings. Conversely, CAC aims to count objects belonging to classes never seen during training, typically operating in a few-shot setting. In this paper, for the first time, we review advancements in CAC methodologies, categorizing them into three paradigms based on how target object classes can be specified: reference-based, reference-less, and open-world text-guided. Reference-based approaches have set performance benchmarks using exemplar-guided mechanisms. Reference-less methods eliminate exemplar dependency by leveraging inherent image patterns. Finally, open-world text-guided methods utilize vision-language models, enabling object class descriptions through textual prompts, representing a flexible and appealing solution. We analyze state-of-the-art techniques and we report their results on existing gold standard benchmarks, comparing their performance and identifying and discussing their strengths and limitations. Persistent challenges – such as annotation dependency, scalability, and generalization – are discussed, alongside future directions. We believe this survey serves as a valuable resource for researchers to understand the progressive developments and contributions over time and the current state-of-the-art of CAC, suggesting insights for future directions and challenges to be addressed.
Index Terms:
Object Counting, Class-agnostic Counting, Few-shot Counting, Prompt-based Counting, Deep Learning, Survey
I Introduction
The goal of object counting is to automatically estimate the number of object instances in still images or video frames [1]. This challenging task has emerged as a prominent research focus within the realm of computer vision, owing to its vast array of practical applications and interdisciplinary significance. For instance, crowd counting [2, 3, 4, 5, 6, 7, 8, 9, 10] is crucial in event management and analysis, holding paramount importance from social [11], psychological [12], political [13], and security [14] perspectives. In smart agriculture, object counting is vital for livestock monitoring [15, 16], pest abundance estimation [17], and for assessing seedling emergence rates by counting plants and trees [18, 19, 20]. Counting techniques are also valuable in traffic monitoring [21, 22, 23, 24, 25], inventory management [26], biomedical applications [27, 28, 29, 30], and ecological surveys [31, 32].
Figure 1: Comparison between class-specific and class-agnostic counting. On the left, we report two examples of class-specific counters, which need individually trained networks on extensive labeled datasets for each object type. On the right, we show the new class-agnostic counting setting, where the counter estimates the number of objects belonging to arbitrary classes. Note that, in the latter scenario, training and test object classes are different.
However, while humans can effortlessly count objects from various visual categories, current state-of-the-art AI-based methods generally focus on enumerating the instances of a single or small set of known classes. Indeed, humans possess an intuitive ability to discern what merits counting when confronted with unfamiliar objects without requiring a reference example or prior comprehension of the object type. In essence, this inherent counting ability comprises two primary facets: recognizing what is noteworthy for counting and detecting repetitions of countable elements. In contrast, even the very recent general-purpose AI agents and Multi-modal Large Language Models (MLLMs) such as LLaVA [33], MiniGPT4 [34], and mPLUG-owl [35], perform poorly on counting objects belonging to open-vocabulary classes [36], and most state-of-the-art counting approaches are class-specific, requiring individually trained networks and, consequently, large labeled datasets for each object type [37] (see left part of Fig. 1).
Figure 2: Overview of class-agnostic counting paradigms. We propose a taxonomy to classify existing CAC methodologies: a) Reference-based approaches rely on annotated bounding box exemplars, which serve as visual prototypes for the object classes to be counted; b) Reference-less techniques relax these requirements, enabling models to automatically identify the dominant class(es) to be counted; and c) Open-world Text-guided methodologies allow the use of textual descriptions to specify the object classes to be considered.
Very recently, research about object counting has shifted towards a new trend that aims to consider arbitrary object categories and attempts to alleviate the problem of the annotation burden (see right part of Fig. 1). The first tentative of creating a model able to count any class, i.e., to perform Class-Agnostic Counting (CAC), is represented by the work of Lu and Zisserman that proposed a Generic Matching Network (GMN) [38]. To achieve this goal, they formulated counting as a matching problem, leveraging the inherent self-similarity property of images, as objects to be counted within an image are typically repeated. GMN was pre-trained with tracking video data, and it exploits an adaptation module to adapt it to domains of interest. However, the latter adaptation module needs dozens to hundreds of examples and, without adaptation, does not perform well in novel classes.
The work that has undergone a detour from standard class-specific counting is [37]. Here, the authors treated counting as a few-shot regression task. Specifically, in this new setting, the inputs are an image and some so-called exemplars; the latter are defined as bounding boxes around the objects of interest from the same input image and are used by the network as object prototypes. At inference time, similar to the few-shot classification task convention [39, 40], the classes of the objects to be counted are entirely different from the ones seen during training, i.e., the provided exemplars belong to arbitrarily unseen categories. This makes few-shot counting very different from the typical counting task, where the training and test classes are the same. It is worth noting that this paradigm does not explicitly learn from limited data – e.g., using (unsupervised) domain adaptation [41, 42] or uncertainty quantification [43]. Instead, this strategy involves using some data in a fully supervised fashion to learn a class-agnostic counting model that can later be applied to unseen object categories. In this way, it is a data-efficient methodology with respect to the unseen object categories.
For the first time, this survey explores and analyzes the leading works existing in the literature tackling class-agnostic counting. Although this task is very recent, it has been addressed by several approaches published in top-tier venues, leading to increasingly better results. Furthermore, its formulation has slightly changed and evolved compared to its original version, progressively reducing the human intervention in the overall process, i.e., the need for annotated exemplars, either at training or inference time. In this context, our contributions are fourfold: (i) we propose a taxonomy to categorize CAC methods, bringing clarity to the literature; (ii) we provide an overview of the architectures of 29 CAC approaches; (iii) we present and compare the performance of the examined techniques on gold-standard benchmarks; (iv) we critically discuss the results from multiple perspectives.
Specifically, the proposed taxonomy organizes the surveyed CAC methods into three categories, also illustrated in Fig. 2. The first category follows the previously mentioned reference-based or few-shot paradigm [37], where annotated bounding box exemplars are required during both training and inference. Subsequently, CAC evolved into a formulation that shapes our second category, where manually specifying object categories is no longer necessary; the algorithm automatically identifies the dominant class(es) to be counted. This approach is known in the literature as reference-less or exemplar-free. Finally, the most recent paradigm, referred to as open-world text-guided, prompt-based, or zero-shot, allows the system to take natural language descriptions of object classes as input, going hand in hand with the latest vision-language foundation models, and constitutes our third category. Based on this taxonomy, we present an overview of the architectures of the examined methods and show their performance on established gold-standard benchmarks from the literature. Specifically, we consider the FSC-147 dataset [37], which includes over 6,000 images across 147 object categories, and the CARPK dataset [44], explicitly designed for vehicle counting and used to assess cross-dataset generalization. Lastly, we critically discuss the results, highlighting the strengths and limitations of the considered methodologies, the persistent challenges in the field, and potential directions for future research.
We organize the rest of this work as follows. In Sec. II, we review the most influential surveys related to object counting, emphasizing the scope of the current work and outlining the search strategy. Sec. III presents the existing methods according to the taxonomy introduced earlier. Next, in Sec. IV, we provide an overview of the available datasets. In Sec. V, we present the results along with a critical discussion. Finally, in Sec. VI, we conclude the paper with insights into potential directions for future research.
II Existing Surveys, Scope, and Search Strategy
While there are several survey papers reviewing counting methods present in the computer vision literature, most of them concentrate on specific object categories, with a particular emphasis on crowd counting. The most cited survey is [45]. In this work, Sindagi et al. revised the papers up to 2017 in detail, focusing on crowd counting and considering both traditional methodologies and techniques based on convolutional neural networks. Furthermore, they described the most used crowd counting datasets, presenting and discussing the results obtained against these benchmarks. More recently, Li et al. published a similar, more updated survey in [46], retracing the last 20 years of object counting, mainly focusing on crowd counting with density estimation using convolutional neural networks. A plethora of other surveys or short surveys have been published over the last few years: a non-exhaustive list includes [47, 48, 49, 50, 51, 52, 53, 54, 55, 56]. Even if the authors of these works look from different angles the topic, considering different taxonomies and perspectives, they always focus on single-class object counting.
Conversely, in this survey, we review the works in the literature concerning class-agnostic object counting, which refers to a new, recent, and different formulation of the classic object counting task. Here, the goal is to count objects belonging to arbitrary categories never seen during the training stage, and this new paradigm brings along completely different solutions, including different methodologies, architectures, and benchmarks. To the best of our knowledge, the only other existing survey in the literature that addresses this topic is [57]. However, in [57], CAC is treated marginally, and the authors did not cover all the existing literature but just a few works, also because the task was just born at the time of writing.
To select papers for this survey, we began by identifying the gold-standard benchmarks for the CAC task. We then conducted a comprehensive search on Google Scholar for papers that cited these benchmarks. To ensure the quality and reliability of the sources, we included only peer-reviewed published papers in our selection. We excluded non-peer-reviewed preprints from arXiv, except for those that offered freely available implementations that could guarantee the reproducibility of their results.
III Existing Approaches
In the following, we review the considered methodologies, categorizing them into three sections based on our proposed taxonomy: reference-based, reference-less, and open-world text-guided approaches. Some techniques incorporate additional modules or slight architectural modifications from their original versions, leading to potential overlaps across these groups. However, we classify each method under its primary category while describing its extensions in other relevant sections as needed. Table I summarizes the surveyed techniques, also providing additional details such as the publication venue and year. Furthermore, Figure 3 presents an overview of selected methods that we consider key milestones, highlighting their relevance over time.
III-A Problem Formulation
State-of-the-art performance in standard class-specific object counting is achieved using supervised deep learning methods, which typically follow two main strategies: counting through detection and counting through regression. The detection-based approach requires prior identifying individual object instances using an object detector [24, 25, 16, 10]. In contrast, regression-based methods learn a mapping between image features and the object count, bypassing the need for explicit localization. This can be achieved directly or, more commonly, through estimating a density map, i.e., a real-valued function. The final estimation of the number of object instances is achieved by summing up the pixel values of the latter [1, 19, 31, 32].
Detectors, while conceptually simpler and capable of handling a wide range of object classes due to the abundance of suitable annotated datasets, struggle with occluded objects. This limitation is the primary reason density-based approaches have largely replaced detection-based techniques. Thus, also most class-agnostic counting techniques follow the density-based paradigm, and CAC existing benchmarks reflect this in terms of label availability. Precisely, the gold standard labels needed for the supervised training of density regressors are dots (points), i.e., ground truth corresponds to binary spatial images with 1’s at the objects’ center location, indicating their existence, and 0’s at other locations without the objects. Since it has been proved to be difficult to train a density regressor with a loss directly based on dot annotations, most existing works generate smoothed target density maps by superimposing Gaussian kernels $G_{\sigma}$ centered in these dot-annotated locations and using a fixed or adaptive spread parameter $\sigma$, depending on the typical object size in the dataset [58, 1].
Formally, we assume to have a collection of $N$ annotated images denoted as $\mathcal{X}={(I_{1},P_{1}),\dots,(I_{N},P_{N})}$, where $I_{i}$ is the $i$-th image and $P_{i}\in\mathbb{R}^{2}$ the set of 2D point coordinates roughly localizing the objects to be counted in image $I_{i}$. The goal is to learn a regression model $f_{\theta}$ producing a density map $D^{map}=f_{\theta}(I)\in\mathbb{R}^{H\times W}$. The number of objects $n$ in an image sub-region $S\subseteq I$ is estimated by integrating $D^{map}$ over $S$, i.e., summing up pixel values in the considered region, $n=\sum_{p\in S}D^{map}_{p}$. The loss function to train the regression model is usually represented by per-pixel functions, such as minimizing the mean squared error between the ground truth and the predicted density maps.
In the few-shot – or reference-based – counting setting, for each image $I_{i}$, we also assume to have a set of $K$ exemplars expressed as bounding box coordinates $B^{E}=\{b_{i}\}_{i=1:k}\in\mathbb{R}^{4}$, denoting single samples depicting the object class to be counted from the same image $I_{i}$ [37]. In the original formulation, $K$ is set to 3 [37]. Furthermore, in few-shot counting, object classes are divided into base classes $C_{b}$ and novel classes $C_{n}$, where $C_{b}$ and $C_{n}$ have no intersection. For each image $I$ from $C_{b}$, both ground truth density maps and exemplars are provided. In contrast, only exemplars are available for images belonging to $C_{n}$, and the goal is to count objects from $C_{n}$ by leveraging the generalization knowledge from $C_{b}$. Instead, in the reference-less and text-guided zero-shot formulations, exemplar bounding boxes are not used. In the reference-less case, the algorithm automatically selects the dominant class(es) to count by analyzing the repetition of object instances. Rather, in the text-guided zero-shot, for each image $I_{i}$, we also assume to have a textual description $T_{i}$ of the object class to be counted.
III-B Reference-based Class-agnostic Counting
TABLE I: List of the considered methods. We summarize the methodologies discussed in this survey, along with some relevant associated details. The methods are categorized based on our proposed taxonomy – reference-based, reference-less, and open-world text-guided.
Method Venue Year Code Point-level Supervision Test-time Adaptation
Reference-based
FamNet [37] CVPR 2021
CFOCNet [59] WACV 2021
VCN [60] CVPRW 2022
RCAC [61] ECCV 2022
BMNet+ [62] CVPR 2022
SPDNC [63] BMVC 2022
MACnet [64] Pattern Recognit. Lett. 2023
SAFECount [65] WACV 2023
LOCA [66] ICCV 2023
ConCoNet [67] Pattern Recognit. Lett. 2023
SSD [68] IJCAI 2024
SATCount [69] Neural Networks 2024
CSTrans [70] Pattern Recognition 2024
CountVers [71] Knowledge-Based Systems 2024
CACViT [72] AAAI 2024
DAVE [73] CVPR 2024
TFPOC [74] WACV 2024
UPC [75] AAAI 2024
PseCo [76] CVPR 2024
Reference-less
RepRPN-Counter [77] ACCV 2022 N/A
RCC [78] arXiv 2022 N/A
CounTR [79] BMVC 2022 N/A
GCNet [80] Pattern Recognition 2024 N/A
Open-world Text-guided
ZSC [81] CVPR 2023
CounTX [82] BMVC 2023
CLIP-Count [83] ACM MM 2023
VLCounter [84] AAAI 2024
GroundingREC [85] CVPR 2024
VA-Count [86] ECCV 2024
Figure 3: Overview of selected CAC methodologies over time. We highlight key milestone methods along a temporal line. Colors represent their category based on our taxonomy: red for reference-based, green for reference-less, and blue for open-world text-guided CAC approaches.
FamNet
The seminal work about reference-based object counting is represented by [37]. Here, Ranjan et al. formulated object counting as a few-shot regression task, defining the primary goal, the first benchmark, and the pioneering methodology. Specifically, as already mentioned in previous sections and illustrated in Fig. 4, they defined the task objective as inferring the number of objects belonging to novel classes unseen during training, given an image and a few exemplar objects. Since the task was novel, the authors introduced the first dataset tailored for it, a collection of images containing objects belonging to 147 categories named FSC-147; labels comprise dots for the training stage, and exemplar bounding boxes exploited both in the supervised learning and at inference time. See also Sec. IV for more details. Finally, the authors proposed the Few Shot Adaptation and Matching Network (FamNet), a novel architecture designed to address this task. Specifically, it includes two main modules: (i) a multi-scale feature extraction module and (ii) a density prediction module. The first consists of the first four blocks picked up from an ImageNet pre-trained ResNet-50 backbone [87], frozen during training. The convolutional feature maps produced by the third and fourth blocks provide the representation of the images $I_{i}$; furthermore, they are also exploited to obtain the multi-scale features of the exemplar images obtained from $B^{E}$ by performing ROI pooling [88] at different scales. Rather, the density prediction module consists of some stacked convolution blocks and upsampling layers; the last one is a $1\times 1$ convolutional layer in charge of predicting the final $2D$ density map $D^{map}$ having the same size as the input image. The input of this second module does not correspond to the features produced by the feature extractor, as usual in the class-specific counting task. Instead, the authors computed the convolution operation between the scaled exemplar features and the image features, obtaining multiple correlation maps, one for each scale, that are then concatenated and fed to the density predictor. The reason behind this is making the density regressor agnostic to the visual categories. During the learning stage, FamNet is trained by minimizing the mean squared error (MSE) between the predicted and the ground truth density map. Finally, this work also proposed a sort of fine-tuning that exploits test data to improve performance further. Specifically, it introduced an adaptation loss that exploits the locations of the exemplars included in the test set to be used only at inference time. It comprises two loss components: (i) the Min-Count loss derived from the intuition that the pixel sum of a predicted density map region corresponding to a bounding box defining an exemplar must be $\geq 1$, and (ii) the perturbation loss that essentially ground to the idea that density values around the location of exemplars in the predicted density map should ideally look like a Gaussian, similar to correlation filters in tracking algorithms [89, 90, 91].
CFOCNet
Roughly simultaneously, Yang et al. introduced a similar solution [59] named Class-agnostic Few-shot Object Counting Network (CFOCNet). The architecture follows the same scheme as FamNet, comprising an encoder and a decoder. The former has two streams for feature extraction, one for the query image and one for the exemplars (here referred to as reference images). It is based on the first three blocks of an ImageNet pre-trained ResNet-50 model [87]. The features concerning the exemplars are matched by computing the convolution operation with the output of a self-attention mechanism [92] applied to the features relating to the query image. These resulting matching maps are finally fused by exploiting a scale-aware fusing mechanism, and the output is fed into the decoder to compute the density map. The decoder is similar to the one exploited in FamNet – some stacked convolution blocks and upsampling layers – and it is supervised by standard MSE loss as well as by SSIM loss [93] to catch local pattern consistency [94].
VCN
Some of the same authors of [37] proposed an extension of FamNet in [60]. Here, they addressed the limitations due to the narrow size of the dataset on which the supervised learning of FamNet was bounded, and they proposed the Vicinal Counting Network (VCN) [60]. The core of VCN is the same as FamNet with some slight changes; additionally, during training, the authors exploited a generator aiming at augmenting and synthesizing images from the vicinity of training samples, increasing the training set. Specifically, the generator is fed with an input image $I\in\mathbb{R}^{H\times W\times 3}$ and a noise vector $z\in\mathbb{R}^{m}$, and it produces an augmented version $I_{z}\in\mathbb{R}^{H\times W\times 3}$ of $I$. This augmented image serves as the input to the feature extractor and the subsequent regressor module, together with the original input $I$. The entire system is then trained jointly in a cooperative fashion. However, unlike the min-max game typical of GANs [95], VCN is trained by employing a standard minimization problem. Alongside the MSE loss for the density regressor optimization, the authors proposed a reconstruction loss in charge of forcing the generator to produce samples close to the original input samples and a diversity loss to avoid the generator learning identity mapping from $I$ to $I_{z}$. Finally, the authors used the same test time adaptation loss of FamNet [37].
RCAC
Another work inspired by FamNet is [61]. First, the authors explored the main limitations of FamNet by analyzing failure cases, and they found that counting performance is highly conditioned by the diversity of the considered exemplars, i.e., intra-class diversity. Specifically, exemplar boxes are limited in number – they potentially do not cover all object instances – and their diversity is not guaranteed since they are provided by annotators subjectively. To tackle this issue, Gong et al. introduced the Robust Class-Agnostic Counter (RCAC), designed to be more robust to intra-class diversity. The architecture follows a two-stream scheme. The first branch consists of a feature extractor similar to the one exploited in FamNet, which takes as input RGB images $I_{i}$. Conversely, the second branch consists of a different feature extractor based on the VGG backbone [96], which takes as input edge images obtained by processing $I_{i}$ throughout an edge detector. This second feature extractor is randomly initialized and, different from the RGB feature extractor, optimized during the training. The reason claimed by the authors behind using edge images is that edges are a sort of class-agnostic knowledge, i.e., shape is a more reliable cue across instances than color. Then, like in [37], ROI pooling is exploited for cropping exemplar box features from the feature maps generated by the two branches. Still, in this case, exemplar features are augmented with a novel feature augmentation module. After that, feature correlation is obtained by computing the convolution operation between the latter and the full feature maps, like in previous works, and, finally, the predicted density map is obtained with a density prediction module identical to the one introduced in FamNet. In detail, the feature augmentation module aims to generate more exemplars of different colors, shapes, and scales in the feature space by combining the provided exemplars. From this perspective, this module shares the same goal of the VCN [60], i.e., increasing the size of the training set.
BMNet+
Another notable work that represents a milestone is [62]. Here, Shi et al. argued that two factors played crucial roles in previous works: (i) feature representation of query images and exemplars, and (ii) similarity metric for matching exemplar and image features. Previous methods either exploited a fixed [37, 60] or a learnable [59] feature extractor but employed a similarity metric relying on the fixed inner product [37, 60, 59, 61, 64], which yielded insufficient matching outcomes. Hence, the authors proposed the Bilinear Matching Network (BMNet+) to improve the similarity modeling. Specifically, the authors designed a learnable bilinear similarity loss to supervise the similarity-matching results, thus going beyond the fixed inner product. The authors took inspiration from metric learning, which aims to embed data into a space where similar samples are pulled close while dissimilar ones are pushed away [97]. In more detail, their similarity loss pulls close the features between the exemplar and target instances while pushing away the features between exemplars and background patches. Furthermore, similar to [59], the authors also proposed a self-attention mechanism to represent self-similarity among features.
SPDCN
Lin et al. [63] focused on the robustness of exemplar feature representations, such as in VCN [60] and RCAC [61], and proposed the Scale-Prior Deformable Convolution Network (SPDCN). The authors argued that most previous works did not consider exemplar scale information during feature extraction or did not do it intuitively. Therefore, they proposed taking advantage of size information provided by the exemplars, discarding features inconsistent with given scales, and consequently extracting more meaningful features. Specifically, they considered a scale-prior backbone consisting of the first ten convolutional layers of an ImageNet pre-trained VGG-19 network [96], where some specific layers have been converted to scale-prior deformable convolutions [98, 99] to exploit the critical scale information from the exemplars. Afterward, these semantic features are split into two branches for segmentation and density estimation, respectively. In the former branch, such as in FamNet and other previous works, ROI pooling [88] is exploited to extract feature maps concerning the exemplars; next, these features are averaged into a representation vector, and a whole similarity map is calculated by computing the cosine similarity between this vector and the image feature map. In the density estimation branch, class-agnostic features are calculated using element-wise multiplication like in [37, 60, 59, 61] between the similarity map and the image feature map. Then, they are fed into a decoder that follows the design of the Pixel Shuffle Crowd Counter (PSCC) [100], which employs pixel-shuffling [101] for the upsampling operations, and that is responsible for predicting the final density map. Finally, another novel contribution of this paper is the usage of the generalized loss [102] to optimize the density predictor. Unlike the MSE loss, utilized in most previous works, generalized loss measures the distance between the predicted density map directly against the dot map, and not the ground truth density map generated convolving Gaussian kernels. In this way, the cost function is adjusted accordingly to the given exemplars, and it is not fixed to the object-specific sizes.
Figure 4: Train and inference high-level overview of reference-based CAC approaches relying on density regression. Following the notation of this survey, CAC models are fed with query images $I_{i}$ and $K$ exemplars expressed as bounding box coordinates $B^{E}=\{b_{i}\}_{i=1:k}\in\mathbb{R}^{4}$. Usually, $K=3$. Exemplars belong to several object classes, which are different at training and test time. The model is in charge of (i) computing feature representations from these inputs that should be agnostic from specific object classes, and (ii) predicting density maps $D^{map}$ from these feature representations. Density-based methods are the standard approach for object counting in crowded scenes. Training typically involves minimizing a per-pixel loss between the predicted and ground-truth density maps, with the final count obtained by summing the pixel values of the predicted density maps.
MACnet
The authors in [64] argued that bounding boxes are not the ideal representations for most exemplars as they often capture undesirable background portions that harm the matching process. Therefore, they proposed exploiting segmentation masks as an alternative, introducing the Mask Augmented Counting network (MACnet). In more detail, the proposed architecture consists of three modules. First, an ImageNet [103] pre-trained ResNet-50 [87] is used as a feature extractor to embed both the query and the exemplars into a feature space. Specifically, its input consists of the query image $I$, the exemplars bounding box coordinates $B^{E}$, and their associated masked counterparts $B_{m}^{E}$. However, unlike previous works, the exemplars are not retrieved with bounding boxes but through their extreme points. The reason claimed by the authors is twofold: (i) annotation with extreme points is approximately four times faster than using bounding boxes [104], and (ii) extreme points are exploited to compute the segmentation masks with a pre-trained Deep Extreme Cut network (DEXTR) [105]. It is also worth noting that while previous literature got the exemplars by cropping them in the feature space, here they are retrieved in the pixel space since elements in the extracted feature tensor already capture background information due to the large receptive field of the feature extractors. The second module consists of a residual layer fed with the feature representation of the exemplars and their masked counterparts. Specifically, it mitigates errors during the segmentation mask creation, learning from the residual information instead of immediately discarding the background information. Finally, the third module is responsible for predicting the density map, supervised by the MSE loss. As in previous papers [37, 59], it is fed with a similarity map obtained by convolving the feature map of the query image with the feature map resulting from the residual module as its filter, i.e., by performing the inner product between these two feature maps. However, here, the final count is derived not directly from integrating the density map but by counting the peaks utilizing a peak-finding algorithm. A final observation concerns the usage of a test time adaptation loss like in [37, 60] aiming to encourage the network in distinguishing background from foreground in the provided exemplars.
SAFECount
You et al. introduced SAFECount in [65], which uses Similarity-Aware Feature Enhancement blocks. These blocks compose of two main modules; (i) a Similarity Comparison Module (SCM), and (ii) a Feature Enhancement Module (FEM). In SCM, query and reference features extracted by an ImageNet pre-trained ResNet-18 [87] are compared in a learnable projection space, and a score map is obtained for each reference image. The similarity map is obtained from the score map by normalizing the exemplar dimension and spatial dimension. The FEM uses the similarity map to weight the reference features, highlighting regions in the query image that are similar to the reference images. These weighted features are then fused into the query features, creating an enhanced representation that emphasizes the areas containing objects of interest.
LOCA
Djukic et al. focused on the strength of exemplar feature representations, arguing that, ideally, the prototypes should generalize over the appearance of the selected object category in the image. They introduced the Low-shot Object Counting network with iterative prototype Adaptation (LOCA) [66]. The core of their solution is the object prototype extraction module, which separately analyzes the shape and appearance characteristics of exemplars and iteratively refines them into object prototypes. Specifically, appearance information is injected by initializing the prototypes with image features computed by RoI pooling [88] exemplar bounding boxes over a query image feature map obtained with a pre-trained ResNet50 [87]. However, the pooling operation makes the appearance queries shape-agnostic, since it maps features from different spatial shapes into rectangular queries of the same size. Thus, the authors injected also shape information, by initializing the prototypes with exemplar width and height features. Finally, the shape and appearance queries are converted into object prototypes by an iterative adaptation module using a recursive sequence of cross-attention blocks.
ConCoNet
Soliven et al. introduced the Contrast Counting Network (ConCoNet) in [67]. They proposed to use not only the (positive) exemplars as in previous works but also the negative exemplars, i.e., object instances defining what not to count. Their contribution can be incorporated into other class-agnostic counting models, and, by way of examples, they reported the results obtained using two of the milestones for the reference-based class-agnostic counting task, e.g., FamNet [37] and BMNet+ [62]. Specifically, the class-agnostic similarity maps are computed not only for the query image and positive exemplar pairs but also for the negative ones; the final similarity map inputted to the density regressor is obtained by simply subtracting the negative similarity map from the positive one. Since CAC datasets do not contain negative exemplars, the authors derived them under the assumption that users provide negative examples during inference, thus performing a sort of adaptation at test time; instead, to enable the training, they selected bounding boxes randomly from the query image regions classified as background by both the ground-truth and the predicted density maps.
SSD
Instead, in [68], the authors leveraged distinct similarity distribution characteristics between several parts of exemplar objects. They argued that different parts of an exemplar object generate distinct similarity distribution patterns. For instance, the similarity distribution at the center of the object gradually decreases from the center toward the outer regions, while the distribution at the edges varies depending on the location. To capture this, they proposed the Spatial Similarity Distribution (SSD) network, which produces a 4D similarity tensor. This allows for flexible extraction of point-to-point similarity distribution information between query and exemplar features through convolution operations in 4D space.
SATCount
Differently, Wang et al. proposed a Scale-Aware Transformer-based class-agnostic counting framework named SATCount [69]. Here, to combine the features from exemplars and the query, the authors developed a feature fusion module comprising two sub-modules: the Scale-Aware Module (SAM) and the Similarity Modeling Module (SMM). SAM combines visual and scale information to generate scale-aware visual features of exemplar objects. Meanwhile, SMM uses cross-attention to model the similarity between exemplar and query features, allowing for long-distance feature associations and the integration of global information from both exemplars and queries.
CSTrans
Gao and Huang [70] drew inspiration from FamNet, enhancing its architecture with two new modules: Correlation-guided Self-Activation (CSA) and Local Dependency Transformer (LDT). The resulting approach, named Correlation-guided Self-Activation Transformer (CSTrans), improves the correlation representation between the exemplar patch and query image features. Additionally, it incorporates a density map predictor based on Transformers [106] that is fully learnable and better equipped to model local context dependencies.
CountVers
The authors in [71] focused on a novel correlation mechanism between exemplar-query features, overcoming the limitations of kernel-wise convolution characterizing previous works such as FamNet. Specifically, they proposed CountVers, which integrates both channel-wise and spatial-wise correlation mechanisms for simultaneously counting small and large objects. The innovative perspective of this work relies on two key insights. On the one hand, small objects in query images tend to have features concentrated in specific spatial regions, highlighting the importance of the channel dimension and thus suggesting channel-wise correlation mechanisms for enhancing the counting accuracy of small objects. On the other hand, large objects often span significant spatial areas, resulting in complex spatial semantic features, and thus utilizing spatial-wise correlation mechanisms effectively is a crucial factor for ensuring accurate counting of large objects.
CACViT
Recently, Zhicheng et al. proposed CACViT [72], a framework for CAC that leverages a single pre-trained standard Vision Transformer (ViT). This work highlights the practicality of using a plain ViT for CAC by introducing a novel "extract-and-match" paradigm, which eliminates the need for multiple feature extractors or post-processing for feature matching. Specifically, the authors demonstrated that the attention mechanism of ViT can not only extract features from both the query and exemplar images but also perform feature matching. By concatenating the query and exemplar tokens and inputting them into a basic ViT model, the self-attention process within the ViT is viewed as comprising two distinct groups: one involving self-attention and the other cross-attention. The self-attention component is responsible for extracting features from both the query and exemplar images, while the cross-attention facilitates the matching between them.
DAVE
Pelhan et al. recently introduced a detect-and-verify paradigm (DAVE) [73] that combines the strengths of density-based methods – best suited for crowded scenarios – with those of detection-based techniques, which provide precise object localization. This approach follows a two-stage pipeline. In the first detection stage, DAVE generates a high-recall set of candidate bounding boxes by first estimating the object centers through a location density map and then predicting bounding box parameters using a regression head. The verification stage refines these detections by removing false positives. Each candidate is analyzed based on its appearance features, which are compared to features of the provided exemplars using clustering methods. Detections that do not align with exemplar clusters are classified as outliers and removed. The refined bounding boxes are then used to update the density map, which further improves count estimation accuracy.
TFPOC
Different from the above-described architectures, TFPOC [74] employed a detection-based technique, leveraging the popular Segment Anything Model [107] for instance segmentation that does not require additional training or fine-tuning for class-agnostic counting. Initially, the authors presented a baseline method. This involves generating binary masks for reference objects using input prompts such as points or boxes, computing a similarity map between the image features and reference object features via cosine similarity, and producing masks for all objects in the image using a grid of point prompts. Target objects are identified by calculating similarity scores for each mask and applying a threshold to determine relevance. The final count is obtained by summing the identified target objects. However, this vanilla approach is computationally intensive as it processes all objects in the image and struggles with setting an optimal similarity threshold, which can lead to undercounting or overcounting. To address these limitations, the authors proposed an advanced prior-guided mask generation method, enhancing the segmentation process with three types of priors. The similarity prior uses a binarized similarity map to guide the segmentation model in focusing on target regions. The segment prior prevents redundant processing by maintaining an overall segment map and refining segmentation iteratively. The semantic prior incorporates reference object features to better guide the mask decoder in identifying and segmenting the target objects.
UPC
A token-based framework for unified open-world text-guided counting (UPC) is introduced in [75]. Prompts, which may take the form of boxes, points, or text, are first transformed into a unified representation called a prompt mask. For box and point prompts, this mask is derived directly from the labeled regions or pixels within the image. As in previous methods, image features are extracted using a convolutional neural network (CNN) encoder, and a cross-attention mechanism is applied to determine the similarity between the prompt token and the image features. This similarity guides the creation of density features, which are subsequently decoded by a CNN to produce the final density map. However, in this work, to improve the model’s robustness, the predicted density map is iteratively refined by treating it as a new prompt mask in successive iterations. This iterative process establishes a fixed point, allowing the model to converge to consistent predictions. A custom-designed fixed-point loss function, grounded in implicit differentiation, further stabilizes training while optimizing parameters effectively. An additional distinctive aspect of this approach is the implementation of a contrastive training strategy. This strategy pairs positive samples, which contain objects matching the prompt, with negative samples, which do not. By compelling the model to generate accurate density maps for positive samples and zero-density maps for negative samples, this method significantly reduces errors associated with noisy or ambiguous prompts.
PseCo
[76] As in TFPOC [74], the authors in [76] exploited a detection-based approach relying on the popular Segment Anything Model [107] for instance segmentation. Specifically, they introduced a framework termed Point, Segment and Count (PseCo). Different from TFPOC, PseCo does not use uniform grid points to prompt the Segment Anything Model to segment all objects. Indeed, the authors argued that using grid points may fall short in capturing all objects, particularly in densely crowded scenes, leading to the omission of many small objects and making this approach unsuitable for object counting. Although increasing the density of grid points could help address this limitation, it would result in substantial computational overhead, as a significant number of points would still target background regions. This makes such an approach impractical for real-world use. Thus, in [76] a class-agnostic object localization that estimates a heatmap of all objects was proposed. From this heatmap, object point coordinates are inferred and are used to feed the Segment Anything Model. Following this step, PseCo introduced a second stage that leveraged another foundation model alongside the Segment Anything Model, namely the popular Contrastive Language–Image Pretraining (CLIP) model [108]. Specifically, CLIP is used to classify the regions found in the first stage with respect to the given exemplars. To encourage a discriminative classification among mask proposals, a novel hierarchical knowledge distillation between mask proposals was also proposed.
III-C Reference-less Class-agnostic Counting
RepRPN-Counter
The first reference-less class-agnostic counting method was introduced in [77]. In this paper, Ranjan et al. introduced the Repetitive Region Proposal Network Counter (RepRPN-Counter), which predicts separate density maps for each repetitive object category in a given query image. The model is based on FamNet [37] and includes an upstream module called the Repetitive Region Proposal Network (RepRPN), designed to automatically identify exemplars of repetitive objects in the image, which are then passed to FamNet for further processing. RepRPN takes inspiration from the Region Proposal Network (RPN) of the popular Faster R-CNN object detector [109]. RPN typically calculates proposal boxes and objectness scores, but RepRPN additionally outputs a so-called repetition score, indicating how frequently an object within the proposals appears in the image. These repetition scores are used to select exemplars from the repetitive object categories in the query image, with the highest-scoring proposals chosen as exemplars. However, the original RPN generated region proposals based on anchors – predefined bounding boxes of varying sizes and aspect ratios used as reference points for object detection – which did not cover the entire image, a critical requirement for computing the global repetition score. To address this, the authors incorporated Encoder Self-Attention layers [106], which can integrate information from similar feature vectors across the entire convolutional feature map of the image, enabling efficient computation of the global repetition score at any anchor location. Another challenge the authors faced was training RepRPN, particularly due to the missing labels in class-agnostic counting datasets such as FSC-147 [37]. Specifically, exemplar bounding boxes and dots are provided for only a single object category, but images may contain other unlabeled categories (see also Sec. IV). However, forcing RepRPN-Counter to predict zero counts for these unannotated objects could negatively impact its performance. To address this issue, the authors employed knowledge transfer from teacher networks. They used RepRPN, trained on the MSCOCO dataset [110], as a teacher network to generate target labels for the objectness and repetition scores of proposals that do not overlap with annotated objects in CAC datasets. Instead, to generate target density maps for the unannotated proposals, they leveraged a pre-trained FamNet, capable of predicting density maps for novel object classes based on a single exemplar.
RCC
[78] went a step further by proposing RCC - Reference-less class-agnostic counting. Unlike previous works, it did not use point-level supervision – besides not using any exemplars as in the reference-less setting – thus proposing a weakly-supervised methodology. Specifically, their proposed network directly regresses to a single scalar count prediction for the entire input image, thus not requiring ground-truth density maps for supervised learning. To create a general and informative feature space without labels, the authors implemented a self-supervised knowledge distillation technique based on [111]. They encouraged the learning of meaningful image representations by fostering agreement between a fixed teacher network, which processes large global image crops, and a student network, which analyzes smaller local image crops. This alignment is achieved by minimizing the cross-entropy between the probability distributions produced by both networks.
CounTR
Another milestone concerning reference-less class-agnostic counting is represented by Counting TRansformer - CounTR [79] (pronounced "counter"). In this work, Liu et al. built on the idea that self-similarity serves as a strong prior in visual object counting. They introduced a transformer-based architecture where self-similarity can be explicitly captured by the inherent attention mechanisms, both between input image patches and with the exemplars (if any). Specifically, this architecture features two visual encoders: (i) a Video Transformer (ViT) [112], which maps input image features into a high-dimensional feature space, and (ii) a lightweight CNN, responsible for extracting visual features from the exemplars (if available). The subsequent feature interaction module is responsible for merging information from both encoders. It consists of a series of standard transformer decoder layers, where the image features serve as the Query, and two separate linear projections of the exemplar features, or a learnable special token if no exemplars are present, act as the Key and Value. This architecture aligns well with the self-similarity prior in counting tasks: self-attention captures similarities between image regions, while the cross-attention between Query and Value compares these regions with the exemplars or learns to bypass the CNN branch when using the learnable token, i.e., in the reference-less setting where no exemplars are provided. Additionally, the authors proposed a two-stage training approach, where the transformer-based image encoder is first pre-trained using self-supervision through masked image modeling [113], followed by supervised fine-tuning for the counting task. Finally, another notable contribution tackled the long-tailed issue in existing counting datasets, where only a few images contain a large number of instances. To handle this, they introduced a scalable pipeline for synthesizing training images with large object counts by cropping random-sized square regions from existing images and arranging them together in a collage.
GCNet
Also the authors of [80] introduced a technique that does not use point-level supervision such as [78], named GCNet - Generalized Counting Network. However, differently from [78], GCNet uses a pseudo-Siamese structure to capture pseudo-exemplar cues from the resized raw image. Specifically, the main branch of the network extracts primary 2D feature maps from an input image $I_{i}$. In contrast, the secondary branch is in charge of generating exemplar-focused outputs by exploiting a methodology inspired by vision transformers [112]: (i) a sequence of overlapping patches is extracted using the unfold operation over the 2D feature maps computed by the main branch, (ii) pseudo-exemplar patches are computed using a per-pixel average operation along the patch dimension, ensuring to capture the common attributes across local regions, (iii) pseudo-exemplar patches are divided into 16 sub-patches of size $2\times 2$, and (iv) each sub-patch is flattened into a 1D token and then transformed through linear projection. The output of the main branch is then further processed through an anisotropic encoder to capture discriminative features along three directions in feature space – horizontal, vertical, and basis/channel directions. Finally, these two outputs are fed into a dual-attention condenser. The same authors proposed also a reference-based version of GCNet, which attaches an additional ResNet-50 backbone to the front-end of the original GCNet for processing labeled exemplar patches; the exemplar features are then fed into the same back-end structure of GCNet to predict an auxiliary single count.
Adaptations
Some existing reference-based techniques have been adapted to the reference-less setting. Specifically, some minor modifications have been implemented to the module of LOCA [66] in charge of extracting object prototypes. Moreover, the authors of DAVE [73] made adjustments to their architecture, replacing the location density prediction component with a reference-less version of LOCA. However, the detection stage and most of the verification stage remained unchanged. The only alteration in the verification stage involves the cluster selection process: clusters with a size of at least 45% of the largest cluster are retained as positive detections, while the others are classified as outliers. This adjustment addresses potential cluster fragmentation due to the lack of exemplars that define appearance similarity levels. Lastly, the authors of [77] proposed a modified version of FamNet, incorporating a standard Region Proposal Network (RPN) from the widely used Faster R-CNN object detector [109] for exemplar generation.
III-D Open-world Text-guided Object Counting
ZSC
Xu et al. [81] were the first to propose a open-world text-guided approach. They highlighted that the earlier reference-based paradigm is not well-suited for autonomous systems, as it relies on human intervention – requiring users to provide exemplar inputs for new categories during inference. Although reference-less class-agnostic counting eliminates the need for bounding box exemplars, it still has limitations in this context. Specifically, it does not allow for user interaction and only estimates the count of objects belonging to the dominant class(es) without identifying the class type. Consequently, its use in autonomous systems is limited to scenarios where the target object class aligns with the majority class(es) present among the objects. Thus, [81] introduces a more versatile paradigm called zero-shot counting (ZSC) where users can specify textual descriptions of the object classes they want to count, representing a sort of trade-off between previous class-agnostic counting paradigms – humans are still in the loop but can specify classes with less effort and in a more natural way. In more detail, their proposal is grounded on an effective strategy that selects meaningful patches in the input image containing the target object specified by the class name to serve as exemplars. To achieve this, they introduced a two-stage technique that (i) builds a class prototype in a pre-trained embedding space based on the given class name and selects patches whose embeddings are the $k$-nearest neighbors to that prototype, and (ii) filters out poor exemplars, retaining only the most relevant ones. Concerning the first step, the authors trained a conditional variational autoencoder (VAE) based on the popular Contrastive Language–Image Pretraining (CLIP) model [108] to generate visual exemplar prototypes conditioned with their semantic embedding encoded in the provided category name. The second stage relies instead on the observation that feature maps from meaningful exemplars show subtle differences compared to those generated from poor exemplars. Specifically, meaningful exemplars tend to display repetitive patterns centered around the object areas, while poor exemplars often exhibit irregular patterns throughout the image. To assess the quality of the generated feature maps – and thus the selected input exemplar patches – the authors employed a pre-trained reference-based object counter similar to FamNet [37], training it to predict the counting error associated with the input exemplar. Once this error predictor was trained, the authors utilized it to filter the selected patches, retaining only those with the smallest predicted errors, and to compute the density maps and associated instance counts in the input images for these exemplars.
CounTX
Concurrently, Amini-Naieni et al. [82] proposed an alternative solution that, unlike the two-stage pipeline in [81], introduced a single-stage approach called CounTX (pronounced "Count-text"). CounTX can be trained end-to-end and accepts more detailed specifications of target objects for counting, rather than relying solely on class names. Specifically, it extends the reference-less CounTR architecture [79] to the open-world text-guided setting. Like CounTR, it features a transformer-based image encoder, but it utilizes the CLIP vision transformer B-16 [108], pre-trained in a contrastive manner with a transformer-based text encoder using image-text pairs from LAION-2B [114]. In more detail, the image encoder encodes each input image into a spatial map of 512-dimensional feature vectors, while the text encoder processes the class description into a single 512-dimensional feature vector. To combine these features, the authors employed a transformer-based feature interaction module, similar to CounTR, but instead of matching image patches to visual exemplars, in CounTX, it is exploited to compute the similarities between image patches and class descriptions. Finally, the output of the feature interaction module is reshaped to a spatial feature map and fed into a progressive four-block convolution and upsampling operation to produce density maps.
CLIP-Count
CLIP-Count [83] is another method built on the Contrastive Language–Image Pretraining (CLIP) model [108], designed to align text embeddings with dense visual features. In this approach, text embeddings are compared with image patch embeddings. Key features of this method include: (i) the use of an InfoNCE-based contrastive loss [115], commonly employed in contrastive learning to encourage meaningful and discriminative representations by pulling similar data points (positive pairs) closer in the latent space and pushing dissimilar ones (negative pairs) further apart. This loss is used alongside the classic MSE loss, applied to binary ground-truth masks that distinguish between positive and negative patch regions; and (ii) a hierarchical patch-text interaction module that propagates semantic information across different resolution levels of visual features. For density map generation, the decoder employs a conventional design based on convolutional and upsampling layers.
VLCounter
In [84], the authors incorporated three modules to efficiently fine-tune CLIP for the counting task and to exploit intermediate features across different encoding layers of CLIP in the decoding stage. In particular: (i) a semantic-conditioned prompt-tuning module is responsible for fine-tuning CLIP efficiently, using conditioning via semantic embeddings to generate patch embeddings that highlight the region of interest, rather than relying on simple learnable prompts; (ii) a learnable affine transformation module adjusts the semantic-patch similarity map produced by the previous module to better suit the counting task; (iii) a segment-aware skip connections module is utilized to capture intermediate features from various encoding layers of CLIP, enhancing the decoder’s generalization ability and providing rich contextual information.
GroundingREC
[85] addressed a key limitation in object counting, where the task is typically confined to the class level and fails to account for fine-grained details within the class. In practical applications, counting often requires contextual or referring human input to target specific objects, such as monitoring traffic flow in different directions or distinguishing between pedestrians and vehicles that are stationary or moving at various points of a junction. To tackle this, the authors introduced a novel task called Referring Expression Counting (REC), which focuses on counting objects with distinct attributes within the same class. To evaluate the REC task, they developed a new dataset named REC-8K, comprising 8011 images and 17,122 referring expressions. The methodology leverages the open-set detector GroundingDino [116], adapted with specific modifications such as replacing bounding boxes with box centers and utilizing the CLS token as a representation of the global semantics of the referring expression (rather than individual text tokens). Additionally, the authors introduced two modules to enhance REC performance: (i) a global-local feature fusion that integrates global image features to infer relational attributes, improving the model’s understanding of the context; (ii) a contrastive learning module that is designed to learn discriminative features associated with attributes by processing multiple referring expressions for different attributes of same-class objects. Besides the REC task, the authors also evaluated their method of class-agnostic counting to show that their approach is generalizable to prior tasks.
VA-Count
In [86], the authors focused on contrasting the challenge of identifying high-quality exemplars effectively. Specifically, they presented the Visual Association-based Zero-shot Object Counting (VA-Count) framework, designed to establish a robust connection between object categories and their visual representations, enabling adaptability across various classes. VA-Count incorporates two key components: the Exemplar Enhancement Module (EEM) and the Noise Suppression Module (NSM), which respectively refine exemplar identification and mitigate adverse effects. The EEM leverages GroundingDino [116], as in [85]. However, here, the authors also introduced a binary filter to refine the candidate exemplars computed by GroundingDino, selecting the ones containing precisely only one object. However, despite the accurate selection of single-object exemplars, errors can arise when exemplars do not belong to the target category, introducing uncertainty in the learning process. The NSM acts as a safeguard by identifying and filtering out such negative exemplars. While the EEM focuses on selecting high-quality exemplars to establish strong visual associations, the NSM uses samples from irrelevant categories to refine these associations further. By employing contrastive learning, the NSM differentiates between positive exemplars (related to the target category) and negative exemplars, reducing the impact of misalignment on the counting task.
Adaptations
Some reference-based techniques have been adapted for the open-world text-guided setting. For example, the authors of DAVE [73] introduced minor adjustments to their architecture, particularly in the cluster selection process during the verification stage. In this approach, the text prompt embedding is derived using the Contrastive Language–Image Pretraining (CLIP) model [108] and compared to the CLIP embeddings of the identified clusters. Cluster embeddings are computed by masking out image regions outside the bounding boxes associated with each cluster and then obtaining the CLIP embedding for the remaining regions. Cosine distances are calculated between the text embedding and each cluster embedding, and clusters with less than 85% of the highest prompt-to-cluster similarity score are classified as outliers. Similarly, the authors of TFPOC [74] enabled textual inputs to define object classes. Their method begins with a coarse similarity map generated using CLIP-Surgery [117], an enhanced variant of CLIP. From this map, reference objects are identified through a series of steps: the similarity map is binarized, the largest connected component (likely containing the target objects) is isolated, and bounding boxes are created around sub-regions within this component for further refinement. These bounding boxes are then used as prompts for Segment Anything Model [107], producing a high-quality similarity map. In another approach, the authors of [75] employed CLIP to generate a prompt mask inspired by MaskCLIP [118]. Cosine distance was utilized to compute the matching score between predefined text features and the corresponding visual features. In PseCo [76], similar to the few-shot setting, the authors used the CLIP text embeddings as classification weights when only known class names.
IV Datasets
There is a lack of publicly available data for the CAC task. Indeed, most existing counting datasets are designed for specific object categories. Notable examples include UCF-QNRF [119], ShanghaiTech [58], and WorldExpo’10 [4] for crowd counting; TRANCOS [120], NDISPark [42, 121], and CARPK [44] for vehicle counting; and VGG Cell [30] for cell estimation. Meanwhile, datasets that include multiple object categories, such as the MSCOCO dataset [110], are unsuitable for counting tasks because they were designed for object detection, typically featuring images with only a small number of object instances. In this section, we outline the limited CAC benchmarks available in the literature. Among them, FSC-147 [37] is considered the gold standard. Additionally, although originally intended for vehicle counting, CARPK [44] is widely used to evaluate generalization capabilities. Fig. 5 illustrates some samples from these two latter benchmarks.
FSC-147
The FSC-147 dataset comprises 6,135 images spanning 147 object categories, including plants, animals, vehicles, and food. These images were sourced from Flickr, Google, and Bing using open-source image scrapers and then manually inspected and filtered. Annotations were manually created by marking dots over the approximate centroids of each object instance – as usual for the object counting task [1]. Occluded instances were annotated only if at least 10% of the object was visible. Additionally, the dataset includes the coordinates of three bounding boxes per image, representing exemplar instances. Each image contains exemplars belonging to a single category, whose natural language name is stored in a text file. If multiple categories were present in an image, one was chosen arbitrarily. Another characteristic of FSC-147 is the significant variation in object counts per image, ranging from 7 to 3,731, with an average of 56 objects per image. The dataset is divided into three splits: 89 object categories are allocated to the training set, while 29 categories are assigned to the validation and testing sets. These splits consist of 3,659, 1,286, and 1,190 images, respectively.
FSC-133
A revised version of FSC-147, called FSC-133, was introduced in [78]. The authors identified and corrected several errors in FSC-147. Specifically, they found that (i) 11 images from the training set also appeared in the validation or testing sets, (ii) some images were effectively duplicates, exhibiting pixel-wise differences of 0 or near 0 when compared at a $224\times 224$ resolution, and (iii) certain duplicate images had inconsistent ground truth counts, with discrepancies of up to 25%. The data split overlaps stemmed from misclassification due to ambiguous or hierarchical categories in FSC-147, such as ‘kidney beans’ and ‘red beans’ or ‘bread rolls’ and ‘baguette rolls.’ In contrast, count discrepancies were observed in images with challenging object occlusions. To address these issues, FSC-133 merged very similar categories and retained only the images with the most accurate counts. The revised dataset consists of 5,898 images across 133 categories, with splits containing 3,877 images for training, 954 for validation, and 1,067 for testing. However, despite the improvements introduced with FSC-133, subsequent works have largely continued to use FSC-147, which remains the standard for this task.
FSC-147-D
Another enhanced version of FSC-147, called FSC-147-D, was introduced by [82] to better support the open-world text-guided object counting paradigm. In this version, Amini-Naieni et al. replaced the simpler object descriptions with more fine-grained and structured natural language sentences for each image, specifying the objects to be counted with greater precision. For example, the original FSC-147 class name for the image with identifier 3696 was updated from "candy pieces" to "the pink desserts garnished with candy pieces".
OmniCount-191 and MCAC
Moreover, two datasets have addressed a key limitation of FSC-147 – its limited number of images featuring multiple object categories – by including images with multiple categories: OmniCount-191 [122] and MCAC [123]. However, OmniCount-191 is not publicly available, and MCAC has significant drawbacks: it lacks appropriate annotations for open-world text-guided approaches and is synthetic, failing to represent realistic scenarios.
CARPK
Finally, several existing papers have also evaluated their performance on the CARPK dataset [44], a collection of drone-view images of parking lots specifically designed for vehicle counting. This benchmark was first utilized for class-agnostic counting by the work that introduced FSC-147 and has since been used by subsequent works, primarily to test cross-dataset generalization. The dataset comprises 1,448 images, divided into a training set of 989 images and a test set of 459 images, with approximately 90,000 annotated vehicles in total. The annotations feature bounding boxes, and density maps are generated using dots placed at the centers of these bounding boxes. Additionally, some of the bounding boxes are utilized to identify exemplars.
Figure 5: Dataset samples. We report some examples of the CAC gold standard benchmarks. a) The FSC-147 dataset [37] is the standard for class-agnostic counting. It includes more than 6,000 images belonging to 147 object classes. We show some sample images together with the provided labels – bounding boxes localizing the exemplars and density maps. b) The CARPK dataset [44] is a set of drone-view images tailored for vehicle counting that is often exploited for assessing the generalization capabilities of CAC models. We show some sample images together with bounding boxes marking the exemplars.
V Results and Discussion
In this section, we present the results of the evaluated methods, following the same taxonomy of previous sections – reference-based, reference-less, and open-world text-guided approaches. First, we describe the experimental setting and then provide a critical analysis of the outcomes across various points of discussion.
V-A Experimental Setting
Some techniques have been adapted to carry out more than one of these tasks (see also Sec. III); in these cases, we report the results for all of them. We considered the two most commonly used benchmarks: FSC-147 [37] and CARPK [44] (see Table II and Table V-A). For FSC-147, in line with existing literature, we report results for both the validation and test subsets. Differently, the CARPK dataset is used to evaluate the models’ generalization capabilities in counting instances of a specific object class – vehicles. Previous works have explored two different settings: i) training models on FSC-147 without using any data from CARPK or the car category in FSC-147, and ii) fine-tuning models using training data from CARPK, where a set of 12 exemplars is randomly sampled from the training set and used as exemplars across all training and test images.
We show the outcomes in terms of two popular counting metrics, following previous literature, i.e., using the mean absolute error (MAE) and the root mean squared error (RMSE), defined as $\text{MAE}=\frac{1}{N_{T}}\sum_{n=1}^{N_{T}}\left|\tilde{c}_{n}-c_{n}\right|$ and $\text{RMSE}=\sqrt{\frac{1}{N_{T}}\sum_{n=1}^{N_{T}}(\tilde{c}_{n}-c_{n})^{2}}$, where $N_{T}$ is the number of test images, and $\tilde{c}_{n}$ and $c_{n}$ are the ground truth and predicted counts, respectively.
TABLE II: Results on the FSC-147 dataset [37]. We report the performance in terms of MAE and RMSE reached by the methodologies discussed in this survey over the gold standard CAC benchmark – FSC-147. The best results are in bold.
Val Set Test Set
Method MAE $\downarrow$ RMSE $\downarrow$ MAE $\downarrow$ RMSE $\downarrow$
Baseline
Mean 53.38 124.53 47.55 147.67
Median 48.68 129.70 47.73 152.46
Reference-based (3 exemplars)
FamNet [37] 23.75 69.07 22.08 99.54
CFOCNet [59] 21.19 61.41 22.10 112.71
VCN [60] 19.38 60.15 18.17 95.60
RCAC [61] 20.54 60.78 20.21 81.86
BMNet+ [62] 15.74 58.53 14.62 91.83
SPDCN [63] 14.59 49.97 13.51 96.80
MACnet [64] 20.65 - 18.38 -
SAFECount [65] 15.28 47.20 14.32 85.54
LOCA [66] 10.24 32.56 10.79 56.97
ConCoNet (FamNet) [67] 18.77 58.96 18.02 91.63
ConCoNet (BMNet) [67] 15.03 57.94 14.21 91.12
SSD [68] 9.73 29.72 9.58 64.13
SATCount [69] 12.27 45.91 10.77 61.14
CSTrans [70] 18.10 58.45 16.38 93.51
CountVers [71] 11.45 42.29 11.28 90.10
CACViT [72] 10.63 37.95 9.13 48.96
DAVE [73] 8.91 28.08 8.66 32.36
TFPOC [74] - - 19.95 132.16
UPC [75] 16.87 59.45 16.68 105.08
PseCo [76] 15.31 68.34 13.05 112.86
CounTR\$ [79] 13.13 49.83 11.95 91.23
GCNet\$ [80] 19.61 66.22 17.86 106.98
Reference-less
RepRPN-Counter (Top5) [77] 29.24 98.11 26.66 129.11
RCC [78] 17.49 58.81 17.12 104.53
CounTR [79] 17.40 70.33 14.12 108.01
GCNet [80] 19.50 63.13 17.83 102.89
FamNet (Top5)\$\$ [37] 32.15 98.75 32.27 131.46
LOCA\$\$ [66] 17.43 54.96 16.22 103.96
DAVE\$\$ [73] 15.54 52.67 15.14 103.49
Open-world Text-guided
ZSC [81] 26.93 88.63 22.09 115.17
CounTX [82] 17.70 63.61 15.73 106.88
CLIP-Count [83] 18.79 61.18 17.78 106.62
VLCounter [84] 18.06 65.13 17.05 106.16
GroundingREC [85] 10.06 58.62 10.12 107.19
VA-Count [86] 17.87 73.22 17.88 129.31
DAVE [73] 15.48 52.57 14.90 103.42
TFPOC+ [74] 47.21 127 24.79 137.15
UPC [75] 16.92 58.92 16.81 105.83
PseCo [76] 23.90 100.33 16.58 129.77
* re-implemented by [62]; \$ reference-based modification of a reference-less method; \$\$ reference-less modification of a reference-based method; $\dagger$ implemented by [77]; $\ddagger$ prompt-based modification of a reference-based method; $+$ results on the validation set computed by [85].
TABLE III: Results on the CARPK dataset [44]. We report the performance in terms of MAE and RMSE obtained with some of the methodologies discussed in this survey over the CARPK benchmark. Following previous literature, CARPK is exploited for testing cross-dataset generalization over a specific object class – vehicles. The best results are in bold.
Method Fine-Tuned MAE $\downarrow$ RMSE $\downarrow$
Baseline
Mean N/A 65.63 72.26
Median N/A 67.88 74.58
Reference-based (12 exemplars)
FamNet [37] 18.19 33.66
RCAC [61] 13.62 19.08
BMNet+ [62] 5.76 7.83
SPDCN [63] 10.07 14.12
SAFECount [65] 5.33 7.04
CSTrans [70] 5.06 6.53
CountVers [71] 5.19 6.55
CACViT [72] 4.91 6.49
[…]
RCAC [61] 17.98 24.21
BMNet+ [62] 10.44 13.77
SPDCN [63] 18.15 21.61
ConCoNet (FamNet) [67] 23.55 32.41
ConCoNet (BMNet) [67] 16.40 21.54
SAFECount [65] 16.66 24.08
LOCA [66] 9.97 12.51
SATCount [69] 9.51 12.67
CACViT [72] 8.30 11.18
SSD [68] 9.58 12.15
TFPOC [74] 10.97 14.24
UPC [75] 7.83 9.74
Open-world Text-guided
CounTX [82] 11.64 14.85
CLIP-Count [83] 11.96 16.61
VLCounter [84] 6.46 8.68
VA-Count [86] 10.63 13.20
TFPOC [74] 11.01 14.34
UPC [75] 7.62 9.71
$\ddagger$ prompt-based modification of a reference-based method.
V-B Reference-based Results
Table II shows the results achieved by the surveyed reference-based methods on the FSC-147 dataset [37]. We additionally report the results concerning two trivial baselines as in [37], i.e., always output the average object count for training images and always output the median count for the training images. The pioneering FamNet [37] laid the groundwork for this paradigm with an MAE and an RMSE of 22.08 and 99.54 on the test subset. Subsequent works progressively improve the performance, sometimes revising and adapting this foundation model. For instance, VCN [60], RCAC [61], and SPDCN [63] focused on the robustness of exemplar feature representations, improving generalization and highlighting the importance of robust training data diversity. Especially VCN and SPDCN reached an impressive MAE of 18.17 and 13.51, respectively. Differently, MACnet introduced segmentation masks as an alternative to bounding boxes for exemplar representation, demonstrating that finer object delineation can improve model robustness. The subsequent method that has marked a breakthrough is BMNet+, which was one of the first works that improved the matching technique between image and exemplar features, moving away from the naive idea of the fixed inner product between them. SAFECount [65] also followed the direction of improving the image-exemplars matching mechanism and achieved an MAE and RMSE of 14.32 and 85.54, respectively (again, on the test subset). Other methods that stood out are SSD [68], LOCA [66], and CACViT [72] which sensibly lowered both MAE and RMSE. The latter approach proposed a particularly interesting "extract-and-match" paradigm, which basically totally relies on the attention mechanism of Video Transformer [112]. Additionally, also ConCoNet is worth mentioning since it proposed a training strategy potentially applicable to all the other methods that brought the idea of leveraging negative exemplars to better distinguish between target and background regions. However, the model achieving state-of-the-art results on FSC-147 is DAVE [73]. By employing a detect-and-verify strategy, DAVE combines bounding-box detection with clustering-based verification to eliminate outliers effectively. On the leaderboard, DAVE demonstrates the lowest MAE and RMSE values to date (8.66 and 32.36, respectively, on the test subset), showcasing its dominance in the reference-based paradigm.
V-C Reference-less Results
In Table II, we report the results obtained by the surveyed reference-less approaches on the FSC-147 dataset [37]. The first reference-less approach, RepRPN-Counter [77], introduced the concept of repetitive region proposals and reached an MAE of 26.66 and an RMSE of 129.11 on the test subset. These error metrics are significantly higher than those obtained with reference-based approaches, as expected since reference-less techniques do not utilize bounding box exemplar annotations. Subsequent methods have significantly improved performance. CounTR [79], a transformer-based reference-less framework, leverages the self-similarity of image regions and achieves competitive results, with an MAE of 14.12 and an RMSE of 108.01 on the test subset. Another notable contribution is RCC [78]. Besides achieving competitive results (an MAE of 17.12 and an RMSE of 104.53), it employs a weakly supervised learning approach that does not need point-level annotations, thereby representing a particularly interesting approach in data scarcity scenarios. In the same direction, also GCNet [80] does not use dot labels and obtained comparable results to RCC. Overall, there is not a clear winner in this leaderboard. Indeed, on the one hand, the reference-less version of DAVE [73] achieves the best results in the validation split with an MAE and RMSE of 15.54 and 52.67 respectively; on the other hand, CounTR and GCNet shine in the test subset with an MAE of 14.12 and an RMSE of 102.89.
V-D Open-world Text-guided Results
In the last part of Table II, we show the results concerning the considered open-world text-guided approaches, which represent a recent and promising advancement in class-agnostic counting, driven by the progress of large multi-modal models. However, similar to the reference-less case, this flexibility comes at the expense of lower performance compared to reference-based techniques. The first milestone, ZSC [81], obtained an MAE and an RMSE of 22.09 and 115.17 on the test subset. Among the subsequent methods that steadily improved performance, CounTX [82], emerges as a particularly competitive approach, in some way adapting its counterpart CounTR [79] to this new setting. It achieves an impressive MAE of 15.73 and an RMSE of 106.88 on the test subset. However, the best results were contended for by the open-world text-guided variant of DAVE [73] and GroundingREC [85]. On the one hand, the first one excels in the RMSE metric, i.e., is more robust to outliers. On the other hand, the latter prevails on the MAE (10.06 and 10.12 on the validation and test subset, respectively).
V-E Cross-dataset Generalization
Table V-A summarizes the results on the CARPK dataset [44], focusing on cross-dataset generalization among the methods that reported their outcomes on this benchmark. Fine-tuned reference-based approaches demonstrate substantial performance gains, with CACViT [72] leading the pack, achieving an MAE of 4.91 and an RMSE of 6.49. This highlights the effectiveness of incorporating transformers for robust feature extraction and attention mechanisms. BMNet+ [62] and SAFECount [65] also deliver strong results, with MAEs of 5.76 and 5.33, respectively, leveraging advanced similarity modules and data-driven refinement strategies. Non-fine-tuned methods generally show higher error rates, yet recent advancements like LOCA [66], SATCount [69], CACViT [72], and SSD [68] deliver impressive performance, narrowing the gap fine-tuned approaches. Notably, the best-performing non-fine-tuned method is UPC [75], achieving an MAE of 7.84 and an RMSE of 9.74. In contrast, for open-world text-guided methods, VLCounter [84] stands out as the top performer, with an MAE of 6.46 and an RMSE of 8.68, even if UPC still achieved very competitive results (MAE of 7.62 and RMSE of 9.71).
V-F Textual Semantic Understanding
Recently, [124] highlighted notable weaknesses in state-of-the-art open-world text-guided class-agnostic counting methods. Specifically, the authors argued that these models often fail to accurately identify which object class needs to be counted based on the textual description. Instead, they tend to count instances of the predominant class, disregarding the true intent of the prompt. To address these shortcomings, the authors introduced a novel benchmark, including a test suite specifically designed to evaluate these limitations in current evaluation systems. Table IV presents results on the FSC-147 dataset, focusing on their negative-label test. This test evaluates the model’s performance on single-class images by prompting it with textual descriptions referencing absent object classes. [124] introduced also two metrics for the assessment of this test: (i) the normalized mean of negative predictions (NMN), which is the absolute counting error computed by prompting the model with the negative classes, normalized by the ground truth of the positive class, and (ii) the positive class count nearness (PCCN) that mixes positive and negative class predictions, providing an overall quantitative assessment of strong failures in the model. More details can be found in [124]. DAVE [73] performs exceptionally well across nearly all metrics proposed in [124], standing out as the best-performing model in this challenging scenario. Although the authors noted that DAVE occasionally produces catastrophic errors due to sporadic outliers, it remains the top performer overall in this evaluation.
TABLE IV: Textual Semantic Understanding Results. We illustrate the results on the FSC-147 dataset as reported in [124], where the authors introduced a test suite designed to evaluate if open-world test-guided CAC methods are truly able to understand the provided textual prompts. The best results are in bold.
Validation Set Test Set
Method NMN $\downarrow$ PCCN $\uparrow$ NMN $\downarrow$ PCCN $\uparrow$
CounTX [82] 0.87 69.79 0.95 64.51
CLIP-Count [83] 1.24 48.11 1.27 38.13
VLCounter [84] 1.07 62.70 1.15 53.36
TFPOC [74] 0.67 62.62 0.75 66.04
DAVE [73] 0.13 95.90 0.08 97.62
VI Conclusion and Open Challenges
Class-agnostic counting (CAC) has become a pivotal field in object counting, addressing the limitations of traditional class-specific methods. Unlike conventional approaches that require extensive labeled data for each object category, CAC enables models to count objects across arbitrary categories, including those unseen during training. For the first time, this survey reviewed the evolution of CAC methodologies, categorizing them into three paradigms: reference-based, reference-less, and open-world text-guided approaches. Each paradigm brings distinct advantages and trade-offs, reflecting the diverse demands of real-world applications.
Reference-based methods, which rely on annotated exemplars, have established strong performance benchmarks by leveraging exemplar-guided mechanisms. These methods excel in precision and robustness, particularly in structured scenarios. However, their reliance on labeled exemplars limits scalability, making them less suitable for tasks requiring minimal human intervention. Reference-less methods address this limitation by automatically identifying dominant object classes in an image, eliminating the need for exemplars. While these approaches demonstrate greater flexibility and ease of use, their performance still lags behind exemplar-based methods, particularly in complex or cluttered scenes. Further refinements are needed to close this gap. The latest paradigm, open-world text-guided counting, introduces the ability to specify object classes through textual prompts, leveraging advancements in vision-language models. This approach offers unprecedented adaptability and aligns closely with the demands of dynamic and diverse environments. However, as with reference-less methods, this flexibility comes at the cost of reduced accuracy compared to exemplar-based approaches.
Through an in-depth analysis of state-of-the-art methods, including benchmarks such as FSC-147 [37] and CARPK [44], this survey highlights notable advancements and persistent challenges. Across paradigms, a clear ranking of models based on performance metrics does not always emerge. DAVE [73] leads the field in reference-based counting on FSC-147, achieving the lowest MAE (8.66) and RMSE (32.36). Its detect-and-verify strategy sets a benchmark for accuracy and robustness. Among reference-less models, instead, the FSC-147 leaderboard remains inconclusive. On the one hand, DAVE excels, achieving the best result on the validation split. On the other hand, CounTR [79] and GCNet [80] stand out in the test subset. The former showcases the effectiveness of its transformer-based design, while the latter is particularly noteworthy for its weakly supervised methodology, which does not rely on point-level annotations. Similarly, in open-world text-guided counting, FSC-147 lacks a clear winner. While DAVE often tops the leaderboard, demonstrating robustness to outliers, GroundingREC [85] achieves the lowest MAE in both the validation and test splits. Regarding generalization capabilities, the CARPK leaderboard indicates that CACViT [72], UPC [75], and VLCounter [84] are the top-performing approaches across different paradigms and experimental settings. Finally, an analysis of textual semantic understanding – applicable only to open-world text-guided methods – reveals that DAVE remains the best-performing approach.
Open challenges still remain in class-agnostic counting and affect all paradigms. The first limitation to overcome is the lack of publicly available datasets. The gold standard, FSC-147, has several drawbacks. First, it contains a limited number of images and object classes. Furthermore, as pointed out in [124], almost all images contain only a single class of objects. Multi-class images would be a desirable feature for new datasets, as objects belonging to classes different from the one being counted could act as distractors for the counting approach. Finally, the provided textual descriptions are also insufficient, as they consist only of class names, whereas a desirable improvement would be to include more fine-grained class descriptions. The lack of diverse and large-scale datasets also leads to challenges in designing learning approaches suited for data-scarce scenarios. To date, only two methodologies –RCC [78] and GCNet [80] – operate with weak supervision, avoiding the use of dot annotations. This setting is particularly promising, as it could help mitigate the challenge of creating large-scale labeled datasets for the counting task. Other challenges are more specific to open-world text-guided methodologies, which seem to represent the future, especially considering the ubiquity of LMMs that we are witnessing. Specifically, they have been shown to not always correctly interpret the meaning of the prompt. Even though some models appear relatively robust, improvements are needed, and new architectural designs and learning approaches – such as contrastive learning over classes not present in the images – should be considered.

### v2

@@
 Abstract
-Object counting has recently shifted towards class-agnostic counting (CAC), which addresses the challenge of counting objects across arbitrary categories, tackling a critical need in versatile counting systems.
-While humans effortlessly identify and count objects from diverse categories without prior knowledge, most counting methods remain restricted to enumerating instances of known classes, requiring extensive labeled datasets for training, and struggling under open-vocabulary settings.
-Conversely, CAC aims to count objects belonging to classes never seen during training, typically operating in a few-shot setting.
-In this paper, for the first time, we review advancements in CAC methodologies, categorizing them into three paradigms based on how target object classes can be specified: reference-based, reference-less, and open-world text-guided.
-Reference-based approaches have set performance benchmarks using exemplar-guided mechanisms.
+Visual object counting has recently shifted towards class-agnostic counting (CAC), which addresses the challenge of counting objects across arbitrary categories – a crucial capability for flexible and generalizable counting systems.
+Unlike humans, who effortlessly identify and count objects from diverse categories without prior knowledge, most existing counting methods are restricted to enumerating instances of known classes, requiring extensive labeled datasets for training and struggling in open-vocabulary settings.
+In contrast, CAC aims to count objects belonging to classes never seen during training, operating in a few-shot setting.
+In this paper, we present the first comprehensive review of CAC methodologies.
+We propose a taxonomy to categorize CAC approaches into three paradigms based on how target object classes can be specified: reference-based, reference-less, and open-world text-guided.
+Reference-based approaches achieve state-of-the-art performance by relying on exemplar-guided mechanisms.
 Reference-less methods eliminate exemplar dependency by leveraging inherent image patterns.
-Finally, open-world text-guided methods utilize vision-language models, enabling object class descriptions through textual prompts, representing a flexible and appealing solution.
-We analyze state-of-the-art techniques and we report their results on existing gold standard benchmarks, comparing their performance and identifying and discussing their strengths and limitations.
-Persistent challenges – such as annotation dependency, scalability, and generalization – are discussed, alongside future directions.
-We believe this survey serves as a valuable resource for researchers to understand the progressive developments and contributions over time and the current state-of-the-art of CAC, suggesting insights for future directions and challenges to be addressed.
+Finally, open-world text-guided methods use vision-language models, enabling object class descriptions via textual prompts, offering a flexible and promising solution.
+Based on this taxonomy, we provide an overview of the architectures of 29 CAC approaches and report their results on gold-standard benchmarks.
+We compare their performance and discuss their strengths and limitations.
+Specifically, we present results on the FSC-147 dataset, setting a leaderboard using gold-standard metrics, and on the CARPK dataset to assess generalization capabilities.
+Finally, we offer a critical discussion of persistent challenges, such as annotation dependency and generalization, alongside future directions.
+We believe this survey will serve as a valuable resource for researchers, highlighting the progressive developments in CAC and offering insights for future advancements in the field.
 Index Terms:
@@
 VI, we conclude the paper with insights into potential directions for future research.
+Figure 3: Overview of selected CAC methodologies over time.
+We highlight key milestone methods along a temporal line.
+Colors represent their category based on our taxonomy: red for reference-based, green for reference-less, and blue for open-world text-guided CAC approaches.
 II Existing Surveys, Scope, and Search Strategy
@@
 We excluded non-peer-reviewed preprints from arXiv, except for those that offered freely available implementations that could guarantee the reproducibility of their results.
-III Existing Approaches
-In the following, we review the considered methodologies, categorizing them into three sections based on our proposed taxonomy: reference-based, reference-less, and open-world text-guided approaches.
-Some techniques incorporate additional modules or slight architectural modifications from their original versions, leading to potential overlaps across these groups.
-However, we classify each method under its primary category while describing its extensions in other relevant sections as needed.
-Table I summarizes the surveyed techniques, also providing additional details such as the publication venue and year.
-Furthermore, Figure 3 presents an overview of selected methods that we consider key milestones, highlighting their relevance over time.
-III-A Problem Formulation
-State-of-the-art performance in standard class-specific object counting is achieved using supervised deep learning methods, which typically follow two main strategies: counting through detection and counting through regression.
-The detection-based approach requires prior identifying individual object instances using an object detector [24, 25, 16, 10].
-In contrast, regression-based methods learn a mapping between image features and the object count, bypassing the need for explicit localization.
-This can be achieved directly or, more commonly, through estimating a density map, i.e., a real-valued function.
-The final estimation of the number of object instances is achieved by summing up the pixel values of the latter [1, 19, 31, 32].
-Detectors, while conceptually simpler and capable of handling a wide range of object classes due to the abundance of suitable annotated datasets, struggle with occluded objects.
-This limitation is the primary reason density-based approaches have largely replaced detection-based techniques.
-Thus, also most class-agnostic counting techniques follow the density-based paradigm, and CAC existing benchmarks reflect this in terms of label availability.
-Precisely, the gold standard labels needed for the supervised training of density regressors are dots (points), i.e., ground truth corresponds to binary spatial images with 1’s at the objects’ center location, indicating their existence, and 0’s at other locations without the objects.
-Since it has been proved to be difficult to train a density regressor with a loss directly based on dot annotations, most existing works generate smoothed target density maps by superimposing Gaussian kernels $G_{\sigma}$ centered in these dot-annotated locations and using a fixed or adaptive spread parameter $\sigma$, depending on the typical object size in the dataset [58, 1].
-Formally, we assume to have a collection of $N$ annotated images denoted as $\mathcal{X}={(I_{1},P_{1}),\dots,(I_{N},P_{N})}$, where $I_{i}$ is the $i$-th image and $P_{i}\in\mathbb{R}^{2}$ the set of 2D point coordinates roughly localizing the objects to be counted in image $I_{i}$.
-The goal is to learn a regression model $f_{\theta}$ producing a density map $D^{map}=f_{\theta}(I)\in\mathbb{R}^{H\times W}$.
-The number of objects $n$ in an image sub-region $S\subseteq I$ is estimated by integrating $D^{map}$ over $S$, i.e., summing up pixel values in the considered region, $n=\sum_{p\in S}D^{map}_{p}$.
-The loss function to train the regression model is usually represented by per-pixel functions, such as minimizing the mean squared error between the ground truth and the predicted density maps.
-In the few-shot – or reference-based – counting setting, for each image $I_{i}$, we also assume to have a set of $K$ exemplars expressed as bounding box coordinates $B^{E}=\{b_{i}\}_{i=1:k}\in\mathbb{R}^{4}$, denoting single samples depicting the object class to be counted from the same image $I_{i}$ [37].
-In the original formulation, $K$ is set to 3 [37].
-Furthermore, in few-shot counting, object classes are divided into base classes $C_{b}$ and novel classes $C_{n}$, where $C_{b}$ and $C_{n}$ have no intersection.
-For each image $I$ from $C_{b}$, both ground truth density maps and exemplars are provided.
-In contrast, only exemplars are available for images belonging to $C_{n}$, and the goal is to count objects from $C_{n}$ by leveraging the generalization knowledge from $C_{b}$.
-Instead, in the reference-less and text-guided zero-shot formulations, exemplar bounding boxes are not used.
-In the reference-less case, the algorithm automatically selects the dominant class(es) to count by analyzing the repetition of object instances.
-Rather, in the text-guided zero-shot, for each image $I_{i}$, we also assume to have a textual description $T_{i}$ of the object class to be counted.
-III-B Reference-based Class-agnostic Counting
 TABLE I: List of the considered methods.
@@
 VA-Count [85] ECCV 2024
-Figure 3: Overview of selected CAC methodologies over time.
-We highlight key milestone methods along a temporal line.
-Colors represent their category based on our taxonomy: red for reference-based, green for reference-less, and blue for open-world text-guided CAC approaches.
+III Existing Approaches
+In the following, we review the considered methodologies, categorizing them into three sections based on our proposed taxonomy: reference-based, reference-less, and open-world text-guided approaches.
+Some techniques incorporate additional modules or slight architectural modifications from their original versions, leading to potential overlaps across these groups.
+However, we classify each method under its primary category while describing its extensions in other relevant sections as needed.
+Table I summarizes the surveyed techniques, also providing additional details such as the publication venue and year.
+Furthermore, Figure 3 presents an overview of selected methods that we consider key milestones, highlighting their relevance over time.
+III-A Problem Formulation
+State-of-the-art performance in standard class-specific object counting is achieved using supervised deep learning methods, which typically follow two main strategies: counting through detection and counting through regression.
+The detection-based approach requires prior identifying individual object instances using an object detector [24, 25, 16, 10].
+In contrast, regression-based methods learn a mapping between image features and the object count, bypassing the need for explicit localization.
+This can be achieved directly or, more commonly, through estimating a density map, i.e., a real-valued function.
+The final estimation of the number of object instances is achieved by summing up the pixel values of the latter [1, 19, 31, 32].
+Detectors, while conceptually simpler and capable of handling a wide range of object classes due to the abundance of suitable annotated datasets, struggle with occluded objects.
+This limitation is the primary reason density-based approaches have largely replaced detection-based techniques.
+Thus, also most class-agnostic counting techniques follow the density-based paradigm, and CAC existing benchmarks reflect this in terms of label availability.
+Precisely, the gold standard labels needed for the supervised training of density regressors are dots (points), i.e., ground truth corresponds to binary spatial images with 1’s at the objects’ center location, indicating their existence, and 0’s at other locations without the objects.
+Since it has been proved to be difficult to train a density regressor with a loss directly based on dot annotations, most existing works generate smoothed target density maps by superimposing Gaussian kernels $G_{\sigma}$ centered in these dot-annotated locations and using a fixed or adaptive spread parameter $\sigma$, depending on the typical object size in the dataset [86, 1].
+Formally, we assume to have a collection of $N$ annotated images denoted as $\mathcal{X}={(I_{1},P_{1}),\dots,(I_{N},P_{N})}$, where $I_{i}$ is the $i$-th image and $P_{i}\in\mathbb{R}^{2}$ the set of 2D point coordinates roughly localizing the objects to be counted in image $I_{i}$.
+The goal is to learn a regression model $f_{\theta}$ producing a density map $D^{map}=f_{\theta}(I)\in\mathbb{R}^{H\times W}$.
+The number of objects $n$ in an image sub-region $S\subseteq I$ is estimated by integrating $D^{map}$ over $S$, i.e., summing up pixel values in the considered region, $n=\sum_{p\in S}D^{map}_{p}$.
+The loss function to train the regression model is usually represented by per-pixel functions, such as minimizing the mean squared error between the ground truth and the predicted density maps.
+In the few-shot – or reference-based – counting setting, for each image $I_{i}$, we also assume to have a set of $K$ exemplars expressed as bounding box coordinates $B^{E}=\{b_{i}\}_{i=1:k}\in\mathbb{R}^{4}$, denoting single samples depicting the object class to be counted from the same image $I_{i}$ [37].
+In the original formulation, $K$ is set to 3 [37].
+Furthermore, in few-shot counting, object classes are divided into base classes $C_{b}$ and novel classes $C_{n}$, where $C_{b}$ and $C_{n}$ have no intersection.
+For each image $I$ from $C_{b}$, both ground truth density maps and exemplars are provided.
+In contrast, only exemplars are available for images belonging to $C_{n}$, and the goal is to count objects from $C_{n}$ by leveraging the generalization knowledge from $C_{b}$.
+Instead, in the reference-less and text-guided zero-shot formulations, exemplar bounding boxes are not used.
+In the reference-less case, the algorithm automatically selects the dominant class(es) to count by analyzing the repetition of object instances.
+Rather, in the text-guided zero-shot, for each image $I_{i}$, we also assume to have a textual description $T_{i}$ of the object class to be counted.
+III-B Reference-based Class-agnostic Counting
 FamNet